How many ml of 1.00M HCl do I need to add to 10.0ml of 1.00M NaOH to get a solution with pH = 1? (Please show your work.)
1
Expert's answer
2020-10-16T08:55:40-0400

СHCl=1,00 M

VNaOH=10.0 ml, CNaOH=1.00 M


HCl + NaOH = NaCl + H2O

H+ + OH- = H2O


СNaOH = nNaOH/VNaOH(L)

nNaOH = СNaOH * VNaOH(L) = 1,00 M * 0, 010 L = 1*10-2 moll


nOH-=nH+


VHCl = nHClHCl = 1*10-2 moll / 1,00 M = 0,01 L = 10 ml


You need 10 ml of 1,00 M HCl